### v1

# UNetMamba: Efficient UNet-Like Mamba for Semantic Segmentation of High-Resolution Remote Sensing Images

Abstract
The semantic segmentation of high-resolution remote sensing images plays a crucial role in downstream applications such as urban planning and disaster assessment. However, existing Transformer-based methods suffer from the constraint between accuracy and efficiency. To overcome this dilemma, we propose UNetMamba, a novel Mamba-based semantic segmentation model. It incorporates a Mamba Segmentation Decoder (MSD) that can efficiently decode the complex information within high-resolution images, and a Local Supervision Module (LSM), which is train-only but can significantly enhance the perception of local contents. Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with the mIoU increased by 0.87% on LoveDA and 0.36% on ISPRS Vaihingen, while achieving high efficiency through light weight, low memory footprint and low computational cost. The source code will soon be publicly available at https://github.com/EnzeZhu2001/UNetMamba.
Index Terms:
Remote sensing, semantic segmentation, mamba
publicationid: pubid:
I Introduction
Driven by advancement in sensor technology, high-resolution remote sensing images have become increasingly accessible. Rich contents not only make them vital in semantic segmentation for downstream applications, but also bring great challenges in accuracy and efficiency due to high information complexity.
The CNN-based method UNet[1] is a milestone in semantic segmentation, which has established the foundational architecture for subsequent works: an encoder-decoder framework with skip connections. With the success of Transformer[2] in computer vision, Wang et al.[3] introduce Transformer into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images. Consequently, numerous subsequent works have focused on innovating Transformer-based methods[4, 5, 6, 7]. Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and large number of parameters[8] have inevitably limited the application efficiency in high-resolution images.
Recently, Mamba[9] has been proposed to address the quadratic computational complexity of Transformer[2], reducing the complexity to linear one while maintaining competitive long-distance relationship modeling capability[8]. Subsequently, Mamba-based visual foundation models[10, 11] have been designed and further applied to several visual tasks. In the field of semantic segmentation, a series of Mamba-based UNets[12, 13, 14] have achieved outstanding performance in medical image segmentation, showing the potential of visual Mamba in segmentation. For optical remote sensing image interpretation, the 2-D selective scanning mechanism of VMamba[10] has demonstrated competitive efficiency, with RSMamba[15] for classification, CDMamba[16] for change detection, and RS3Mamba[17] for semantic segmentation.
To overcome the dilemma between accuracy and efficiency and then achieve efficient semantic segmentation of high-resolution remote sensing images, we propose a novel UNet-like model named UNetMamba. It consists of three main components: an encoder using ResT[18] backbone, a mamba segmentation decoder (MSD) for efficiently decoding, and a train-only local supervision module (LSM) to enhance the perception of local semantic information. The contributions of our research are summarized as follows:
1) We devise MSD, a plug-and-play Mamba-based decoder for efficiently decoding semantic information from multi-scale feature maps. Specifically, VSS blocks of VMamba[10] are transferred to the decoding side, significantly reducing the parameter count while leveraging the long-distance relationship modeling capability to decode complex information.
2) We also conceive a CNN-based LSM to compensate for insufficient perception of local semantic information, namely, small land-cover elements. With two different scales of convolutions and an auxiliary loss function, we enhance the perception of local semantic information through supervised training. The train-only design also saves on inference cost, further ensuring the application efficiency.
3) Based on MSD and LSM, we propose UNetMamba, a UNet-like model for efficient semantic segmentation of high-resolution remote sensing images. It not only has the advantages of being lightweight and having low memory and computation cost, but also achieves state-of-the-art (SOTA) performance on two well-known high-resolution remote sensing imagery datasets.
Figure 1: Framework of the proposed UNetMamba. (a) Overall architecture of UNetMamba. (b) Visual State Space (VSS) block of Mamba Segmentation Decoder (MSD). (c) Local Supervision Module (LSM) Block of LSM.
II Methodology
As shown in Fig. 1(a), the overall architecture of UNetMamba is constructed under a U-shape framework [1], where the ResT backbone[18] is selected as encoder while decoder is the proposed MSD. Moreover, the LSM is designed to enhance the perception of local semantic information.
II-A Mamba Segmentation Decoder
In previous studies[12, 13, 14, 17], the benefits of employing VMamba[10] for remote sensing semantic segmentation have been established. However, models that utilize a pre-trained VMamba backbone while encoding tend to be overweight for our goal of high efficiency. Therefore, we incorporate the basic unit of VMamba named VSS block[10] into the decoding side only.
The structure of a VSS block is depicted in Fig. 1(b). The 2-D feature map $F$, which has undergone patch expansion, firstly goes through layer normalization and is then linearly embedded:
$F^{\prime}=\text{LinearEmbed}\left(\text{LayerNorm}\left(\text{PatchExp}\left(F\right)\right)\right)$ (1)
Subsequently, the output splits into two information flows. One flow undergoes a 3 × 3 depth-wise convolution followed by a Silu activation function:
$F^{\prime\prime}=\text{SiLU}\left(\text{DWConv}\left(F^{\prime}\right)\right)$ (2)
The SS2D (2-D Selective Scan) module then scans $F^{\prime\prime}$ in four different directions, decoding the semantic information under a global receptive field and linear complexity:
$\begin{split}F_{v}&=\text{ScanExp}\left(F^{\prime\prime},v\right),\quad v\in\left\{1,2,3,4\right\}\\ \bar{F}_{v}&=\text{S6}\left(F_{v}\right),\quad v\in\left\{1,2,3,4\right\}\\ \bar{F}&=\text{ScanMerge}\left(\bar{F}_{1},\bar{F}_{2},\bar{F}_{3},\bar{F}_{4}\right)\end{split}$ (3)
where the expansion and merging operation correspond to VMamba[10], while the S6 operation denotes the selective scan space state sequential model of Mamba[9]. The output of SS2D subsequently undergoes layer normalization and an element-wise multiplication with the other flow, which has also gone through the same linear embedding and Silu activation. Finally, the VSS block outputs through a linear layer with residual connection.
Based on aforementioned operations, the MSD decodes multi-scale feature maps at four different stages, ultimately outputting the semantic segmentation results through a 1 × 1 convolution head.
TABLE I: Quantitative Comparison Results on the LoveDA Dataset at a size of 1024 × 1024 pixels. The best values are in bold.
Method Backbone Background Building Road Water Barren Forest Agriculture mIoU(%)
BANet[3] ResT-Lite 41.92 54.13 52.39 76.18 11.74 45.07 51.71 47.59
ABCNet[4] ResNet-18 40.20 55.01 53.63 76.29 18.43 44.87 59.78 49.74
MANet[5] ResNet-50 44.38 56.20 56.03 78.41 17.06 45.81 58.17 50.87
DC-Swin[6] Swin-T 40.75 57.81 56.55 79.59 16.58 46.49 57.76 50.79
UNetFormer[7] ResNet-18 45.47 58.75 56.53 80.33 18.90 45.46 61.92 52.48
RS3Mamba[17] R18-VMamba-T 41.60 58.23 54.03 77.34 17.97 43.81 61.37 50.62
UNetMamba(ours) ResT-Lite 47.08 59.16 56.74 81.37 18.15 46.61 64.31 53.35
TABLE II: Quantitative Comparison Results on the ISPRS Vaihingen Dataset at a size of 1024 × 1024 pixels.
The metrics are measured on a single NVIDIA RTX 4090. The best values are in bold.
Method Backbone Param(M) Memo(MB) FLOPs(G) mF1(%) mIoU(%) OA(%)
BANet[3] ResT-Lite 12.73 194.61 85.43 90.30 82.43 91.93
ABCNet[4] ResNet-18 13.67 206.74 62.62 90.36 82.49 91.94
MANet[5] ResNet-50 35.86 547.87 216.82 90.65 83.04 92.26
DC-Swin[6] Swin-T 45.63 694.62 190.04 90.66 83.07 92.28
UNetFormer[7] ResNet-18 11.72 179.34 47.03 90.56 82.89 92.20
RS3Mamba [17] R18-VMamba-T 43.32 662.21 157.89 90.71 83.06 92.23
UNetMamba(ours) ResT-Lite 14.76 225.71 100.52 90.91 83.43 92.46
II-B Local Supervision Module
Due to the large receptive field of VSS blocks, the MSD tends to partially overlook local semantic information while decoding . However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation. Therefore, the LSM is proposed to enhance the perception of local details, as depicted in Fig. 1(c). Considering the relatively smaller scale of land-covers in high-resolution remote sensing scenarios, we adopt two parallel convolution branches in LSM with the kernel size set as 3 and 1, respectively. Each branch is followed by a batch normalization layer with a ReLU6 activation function:
$\tilde{F^{\prime}}_{i}=\text{ReLU6}\left(\text{BatchNorm}\left(\text{Conv}_{i}\left(F_{MSD}\right)\right)\right)$ (4)
where $i\in\left\{1,3\right\}$ denotes the kernel size, while $F_{MSD}$ denotes the output feature of MSD at the corresponding stage. The two branches are then merged by addition:
$\tilde{F^{\prime}}=\tilde{F^{\prime}}_{1}\oplus\tilde{F^{\prime}}_{3}$ (5)
Eventually, through a dropout layer, a 1 × 1 convolution consistent with the segmentation head in MSD, and an upsampling layer to the original image size, LSM obtains $\tilde{F}\in\mathbb{R}^{H\times W\times 3}$:
$\tilde{F}=\text{Upsample}\left(\text{Conv}_{1}\left(\text{Drop}\left(\tilde{F^{\prime}}\right)\right)\right)$ (6)
During the training process, we add LSM blocks with different input channels to the decoder at stage 2-4 and then attain the segmentation result by summing up, which is further used in calculating the auxiliary loss function.
II-C Loss Functions
To ensure the holistic performance of UNetMamba while achieving the supervisory role of LSM, loss functions utilized in training are divided into a principal one and an auxiliary one. Following prior works[3, 4, 5, 6, 7], we use a dice loss $L_{dice}$ and a cross-entropy loss $L_{ce}$ to jointly form our principal segmentation loss function:
$L_{dice}=1-\frac{2}{N}\sum_{n=1}^{N}{\sum_{k=1}^{K}{\frac{\hat{y}_{k}^{\left(n\right)}y_{k}^{\left(n\right)}}{\hat{y}_{k}^{\left(n\right)}+y_{k}^{\left(n\right)}}}}$ (7)
$L_{ce}=-\frac{1}{N}\sum_{n=1}^{N}{\sum_{k=1}^{K}{\hat{y}_{k}^{\left(n\right)}\log y_{k}^{\left(n\right)}}}$ (8)
where $N$ denotes the number of samples while $K$ denotes the number of categories. $\hat{y}_{k}^{\left(n\right)}$ is the $k$-th element in one-hot encoding of the true semantic labels for sample $n$. And $y_{k}^{\left(n\right)}$ represents the confidence of sample $n$ belonging to category $k$. As to the auxiliary loss function, cross-entropy loss $L_{ce}$ is selected[7]. Therefore, the overall loss function is formulated as follows:
$L=(L_{dice}+L_{ce})+\alpha L_{ce}$ (9)
where the weight factor $\alpha$ is set as 0.4.
Figure 2: Qualitative comparisons on the LoveDA dataset at size 1024 × 1024. (a) Origin, (b) GT, (c) our proposed UNetMamba, (d) BANet, (e) ABCNet, (f) MANet, (g) DC-Swin, (h) UNetFormer, and (i) RS3Mamba.
Figure 3: Qualitative comparisons on the ISPRS Vaihingen dataset at size 1024 × 1024. (a) Origin, (b) GT, (c) our proposed UNetMamba, (d) BANet, (e) ABCNet, (f) MANet, (g) DC-Swin, (h) UNetFormer, and (i) RS3Mamba.
III Experimental results
III-A Datasets
1) LoveDA: The LoveDA dataset[19] contains 5987 high-resolution remote sensing images at the same size of 1024 × 1024 pixels and includes 7 land-cover categories (background, building, road, water, barren, forest and agriculture). The dataset encompasses two scenes (urban and rural) collected from three Chinese cities (Nanjing, Changzhou and Wuhan), which brings considerable challenges due to multi-scale land-covers and complex contents. In our experiments, following the official settings[19], 2522 images were utilized for training, 1669 images for validation and 1796 images for testing without further cropping.
2) ISPRS Vaihingen: The ISPRS Vaihingen dataset contains 33 TOP (True OrthoPhoto) images, each with very high spatial resolution at an average size of 2494 × 2064 pixels. This widely-used remote sensing dataset involves 6 land-cover categories (impervious surface, low vegetation, building, tree, car and background). In the experiments, we randomly selected 9 images (area ID: 1, 5, 11, 15, 21, 26, 30, 33, 38) for testing, with the remaining 24 images for training, and cropped all of the images into patches at a size of 1024 × 1024 pixels.
III-B Experimental Settings
All of the experiments were implemented with PyTorch on a single NVIDIA RTX 4090 GPU, and the optimizer was set as AdamW with a 0.0006 learning rate and a 0.00025 weight decay. For both of the datasets, the training epoch was set as 100 while the batch size was 8.
Moreover, the number of model parameters (Param), memory footprint (Memo) and floating point operation count (FLOPs) were selected for efficiency evaluation, while the mean F1-score (mF1), mean intersection over union (mIoU) and overall accuracy (OA) were chosen as accuracy evaluation metrics.
III-C Performance Comparison
We conducted a comprehensive performance comparison on LoveDA and ISPRS Vaihingen datasets. A series of SOTA optical remote sensing imagery semantic segmentation models were selected as our competitors, including Transformer-based models: BANet[3], ABCNet[4], MANet[5], DC-Swin[6], UNetFormer[7], and our baseline, RS3Mamba[17].
1) Comparison on the LoveDA Dataset: As shown in Tab. I, the proposed UNetMamba achieved the best accuracy performance in quantitative comparison, with a significant improvement of 0.87% in mIoU. It is note-worthy that our UNetMamba achieved SOTA in six out of seven categories. In background and agriculture, UNetMamba led the second place by 1.61% and 2.39%, respectively. This distinct advantage indicated that, thanks to MSD, our UNetMamba gained excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition. The outstanding performance of UNetMamba in building, road and other small land-cover categories proved that the train-only LSM successfully enhanced the perception of local contents through supervision and generalization. The qualitative comparison results on LoveDA are illustrated in Fig. 2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate.
2) Comparison on the ISPRS Vaihingen Dataset: The proposed UNetMamba not only achieved SOTA performance on the challenging LoveDA dataset, but also performed noticeably well on the classic Vaihingen dataset. As shown in Tab. II, in terms of accuracy, our UNetMamba achieved improvements of 0.20% in mF1, 0.36% in mIoU, and 0.18% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) and low cost (225.71MB, 100.52G). Fig. 3 illustrates the qualitative comparison results, with the framed areas further demonstrating the competitiveness of UNetMamba in the perception of local semantic details, which is particularly important for the semantic segmentation accuracy in high-resolution remote sensing scenes.
TABLE III: Ablation Study results on the LoveDA and ISPRS Vaihingen Dataset. The metrics are measured by a 1024 × 1024 input on a single NVIDIA RTX 4090, and the best values are in bold.
 MSD LSM Param(M) FLOPs(G) mIoU(%)
LoveDA Vaihingen
Baseline with ResT-Lite 42.14 183.64 52.64 83.21
 ✓ 13.89 100.52 52.61 83.15
 ✓ 14.76 100.52 53.35 83.43
III-D Ablation Studies
To assess the effectiveness of the proposed MSD and LSM in UNetMamba, ablation studies were conducted on both of the datasets, with the results listed in Tab. III. For the sake of credibility, the baseline model, RS3Mamba[17], was updated by replacing its main encoding branch with the ResT[18] backbone consistent with UNetMamba. And after removing its Mamba-based auxiliary branch and upgrading its decoder into our MSD, the model experienced significant reduction in Param and FLOPs, with only minor drop in mIoUs, demonstrating the high efficiency of MSD in decoding complex semantic information. On both datasets, deployment of LSM also led to 0.74% and 0.28% increase in mIoUs, respectively. However, the cost was merely a modest parameter-count increase of 0.87M and a constant computation complexity due to the train-only scheme, which further confirming the high efficiency of LSM in enhancing the perception of local semantic information.
IV Conclusion
In this paper, an efficient semantic segmentation model called UNetMamba is proposed for high-resolution remote sensing images. Considering the multi-scale land-covers and complex information in such images, our UNetMamba features a plug-and-play Mamba-based segmentation decoder MSD for efficient semantic information decoding, and a train-only local supervision module LSM that can efficiently enhance the perception of local semantic details. Extensive experiments conducted on two well-known remote sensing datasets demonstrate that UNetMamba not only outperforms other SOTA models, but also achieves lightweight and low costs. In the future, we will continue to explore the potential of novel linear modeling mechanisms in efficient interpretation of high-resolution remote sensing images.

### v2

@@
-UNetMamba: Efficient UNet-Like Mamba for Semantic Segmentation of High-Resolution Remote Sensing Images
+UNetMamba: An Efficient UNet-Like Mamba for Semantic Segmentation of High-Resolution Remote Sensing Images
 Abstract
-The semantic segmentation of high-resolution remote sensing images plays a crucial role in downstream applications such as urban planning and disaster assessment.
-However, existing Transformer-based methods suffer from the constraint between accuracy and efficiency.
-To overcome this dilemma, we propose UNetMamba, a novel Mamba-based semantic segmentation model.
+Semantic segmentation of high-resolution remote sensing images is vital in downstream applications such as land-cover mapping, urban planning and disaster assessment.
+Existing Transformer-based methods suffer from the constraint between accuracy and efficiency, while the recently proposed Mamba is renowned for being efficient.
+Therefore, to overcome the dilemma, we propose UNetMamba, a UNet-like semantic segmentation model based on Mamba.
 It incorporates a mamba segmentation decoder (MSD) that can efficiently decode the complex information within high-resolution images, and a local supervision module (LSM), which is train-only but can significantly enhance the perception of local contents.
-Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with the mIoU increased by 0.87% on LoveDA and 0.36% on ISPRS Vaihingen, while achieving high efficiency through light weight, low memory footprint and low computational cost.
-The source code will soon be publicly available at https://github.com/EnzeZhu2001/UNetMamba.
+Extensive experiments demonstrate that UNetMamba outperforms the state-of-the-art methods with mIoU increased by 0.87% on LoveDA and 0.36% on ISPRS Vaihingen, while achieving high efficiency through the lightweight design, less memory footprint and reduced computational cost.
+The source code is available at https://github.com/EnzeZhu2001/UNetMamba.
 Index Terms:
@@
 I Introduction
-Driven by advancement in sensor technology, high-resolution remote sensing images have become increasingly accessible.
-Rich contents not only make them vital in semantic segmentation for downstream applications, but also bring great challenges in accuracy and efficiency due to high information complexity.
+Spurred by the progress in advanced imaging techniques, high-resolution remote sensing images have become increasingly accessible, which can provide richer contents than regular images.
+Such intricate contents render them crucial for semantic segmentation in several downstream applications, yet simultaneously present significant challenges regarding accuracy and efficiency due to the high complexity of information.
+As the foundation models of computer vision continuously evolve, research on semantic segmentation of high-resolution remote sensing images has flourished.
 The CNN-based method UNet[1] is a milestone in semantic segmentation, which has established the foundational architecture for subsequent works: an encoder-decoder framework with skip connections.
-With the success of Transformer[2] in computer vision, Wang et al.[3] introduce Transformer into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images.
+With the success of Transformer[2], Wang et al.[3] introduce it into the U-shape framework, significantly improving the semantic segmentation accuracy of high-resolution remote sensing images.
 Consequently, numerous subsequent works have focused on innovating Transformer-based methods[4, 5, 6, 7].
-Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and large number of parameters[8] have inevitably limited the application efficiency in high-resolution images.
-Recently, Mamba[9] has been proposed to address the quadratic computational complexity of Transformer[2], reducing the complexity to linear one while maintaining competitive long-distance relationship modeling capability[8].
-Subsequently, Mamba-based visual foundation models[10, 11] have been designed and further applied to several visual tasks.
-In the field of semantic segmentation, a series of Mamba-based UNets[12, 13, 14] have achieved outstanding performance in medical image segmentation, showing the potential of visual Mamba in segmentation.
-For optical remote sensing image interpretation, the 2-D selective scanning mechanism of VMamba[10] has demonstrated competitive efficiency, with RSMamba[15] for classification, CDMamba[16] for change detection, and RS3Mamba[17] for semantic segmentation.
-To overcome the dilemma between accuracy and efficiency and then achieve efficient semantic segmentation of high-resolution remote sensing images, we propose a novel UNet-like model named UNetMamba.
-It consists of three main components: an encoder using ResT[18] backbone, a mamba segmentation decoder (MSD) for efficiently decoding, and a train-only local supervision module (LSM) to enhance the perception of local semantic information.
+Despite the prosperity of Transformer in remote sensing semantic segmentation, its quadratic computational complexity and high parameter count[8] have inevitably limited the application efficiency in high-resolution images.
+Recently, Mamba [9] has become a hot point.
+For efficiency, it successfully achieves a linear complexity, while guaranteeing accuracy through its competitive long-distance dependency modeling capability.
+Subsequently, Mamba-based visual foundation models[10, 11] have been well-designed.
+Benefiting from such models, a series of Mamba-based UNets[12, 13, 14] have achieved impressive accuracy in medical imagery segmentation, showing the potential of visual Mamba in segmentation.
+For remote sensing image interpretation, the 2-D selective scanning of VMamba[10] has also demonstrated promising efficiency, with RSMamba[15] for classification, CDMamba[16] for change detection, and RS3Mamba[17] for semantic segmentation.
+However, these pioneering works have primarily validated the effectiveness of Mamba in their respective tasks, without fully leveraging its dual advantages in accuracy and efficiency through targeted design.
+To overcome the dilemma between accuracy and efficiency and then achieve efficient semantic segmentation of high-resolution remote sensing images, we propose a Mamba-based UNet-like model named UNetMamba.
+It consists of three main components: an encoder using ResT[18] backbone, a mamba segmentation decoder (MSD) for efficiently decoding, and a neatly-designed local supervision module (LSM) to enhance the perception of local semantic information.
 The contributions of our research are summarized as follows:
 1) We devise MSD, a plug-and-play Mamba-based decoder for efficiently decoding semantic information from multi-scale feature maps.
-Specifically, VSS blocks of VMamba[10] are transferred to the decoding side, significantly reducing the parameter count while leveraging the long-distance relationship modeling capability to decode complex information.
-2) We also conceive a CNN-based LSM to compensate for insufficient perception of local semantic information, namely, small land-cover elements.
+Specifically, visual state space (VSS) blocks of VMamba[10] are expressly transferred to the decoding side, significantly reducing the parameter count while leveraging the long-distance modeling capability to accurately decode complex information in a global receptive field.
+2) We also conceive a CNN-based LSM to address the lack of local information perception in MSD.
 With two different scales of convolutions and an auxiliary loss function, we enhance the perception of local semantic information through supervised training.
-The train-only design also saves on inference cost, further ensuring the application efficiency.
+The train-only design also saves on inference costs, further ensuring the application efficiency.
 3) Based on MSD and LSM, we propose UNetMamba, a UNet-like model for efficient semantic segmentation of high-resolution remote sensing images.
-It not only has the advantages of being lightweight and having low memory and computation cost, but also achieves state-of-the-art (SOTA) performance on two well-known high-resolution remote sensing imagery datasets.
+It has the advantages of being lightweight and requiring low costs, while also achieving state-of-the-art (SOTA) performance on two well-known high-resolution remote sensing imagery datasets.
+II Methodology
+As shown in Fig.
+1(a), the overall architecture of UNetMamba is constructed under a U-shape framework [1], where the ResT backbone[18] is selected as encoder while decoder is the proposed MSD.
+Moreover, the LSM is designed to enhance the perception of local semantic information.
 Figure 1: Framework of the proposed UNetMamba.
@@
 (c) Local Supervision Module (LSM) Block of LSM.
-II Methodology
-As shown in Fig.
-1(a), the overall architecture of UNetMamba is constructed under a U-shape framework [1], where the ResT backbone[18] is selected as encoder while decoder is the proposed MSD.
-Moreover, the LSM is designed to enhance the perception of local semantic information.
 II-A Mamba Segmentation Decoder
@@
 However, models that utilize a pre-trained VMamba backbone while encoding tend to be overweight for our goal of high efficiency.
-Therefore, we incorporate the basic unit of VMamba named VSS block[10] into the decoding side only.
-The structure of a VSS block is depicted in Fig. 1(b). The 2-D feature map $F$, which has undergone patch expansion, firstly goes through layer normalization and is then linearly embedded:
+Therefore, we incorporate the basic unit of VMamba, namely, VSS block[10], into the decoding side only.
+The structure of a VSS block is depicted in Fig.
+1(b).
+The 2-D feature map $F$, which has undergone patch expansion, firstly goes through layer normalization and then splits into two information flows.
+The main flow is further linearly embedded as $F^{\prime}$:
 $F^{\prime}=\text{LinearEmbed}\left(\text{LayerNorm}\left(\text{PatchExp}\left(F\right)\right)\right)$ (1)
-Subsequently, the output splits into two information flows. One flow undergoes a 3 × 3 depth-wise convolution followed by a Silu activation function:
+Subsequently, $F^{\prime}$ undergoes a 3 × 3 depth-wise convolution and is then SiLU-activated into $F^{\prime\prime}$:
 $F^{\prime\prime}=\text{SiLU}\left(\text{DWConv}\left(F^{\prime}\right)\right)$ (2)
-The SS2D (2-D Selective Scan) module then scans $F^{\prime\prime}$ in four different directions, decoding the semantic information under a global receptive field and linear complexity:
+The 2-D Selective Scan (SS2D) module thereafter scans $F^{\prime\prime}$ in four different directions, decoding semantic information under a global receptive field and linear complexity:
 $\begin{split}F_{v}&=\text{ScanExp}\left(F^{\prime\prime},v\right),\quad v\in\left\{1,2,3,4\right\}\\ \bar{F}_{v}&=\text{S6}\left(F_{v}\right),\quad v\in\left\{1,2,3,4\right\}\\ \bar{F}&=\text{ScanMerge}\left(\bar{F}_{1},\bar{F}_{2},\bar{F}_{3},\bar{F}_{4}\right)\end{split}$ (3)
 where the expansion and merging operation correspond to VMamba[10], while the S6 operation denotes the selective scan space state sequential model of Mamba[9].
-The output of SS2D subsequently undergoes layer normalization and an element-wise multiplication with the other flow, which has also gone through the same linear embedding and Silu activation.
-Finally, the VSS block outputs through a linear layer with residual connection.
+The $\bar{F}$ then goes through layer normalization and an element-wise multiplication with the bypass flow, which has also undergone the same linear embedding and Silu activation.
+Finally, the VSS block outputs $F_{MSD}$ through a linear layer with residual connection.
 Based on aforementioned operations, the MSD decodes multi-scale feature maps at four different stages, ultimately outputting the semantic segmentation results through a 1 × 1 convolution head.
+II-B Local Supervision Module
+Due to the large receptive field of VSS blocks, MSD tends to partially overlook local semantic information while decoding .
+However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation.
+Therefore, the LSM is proposed to enhance the perception of local details, as depicted in Fig.
+1(c).
 TABLE I: Quantitative Comparison Results on the LoveDA Dataset at a size of 1024 × 1024 pixels. The best values are in bold.
@@
 UNetMamba(ours) ResT-Lite 14.76 225.71 100.52 90.91 83.43 92.46
-II-B Local Supervision Module
-Due to the large receptive field of VSS blocks, the MSD tends to partially overlook local semantic information while decoding .
-However, such details in high-resolution images are key factors to further improve the accuracy of semantic segmentation.
-Therefore, the LSM is proposed to enhance the perception of local details, as depicted in Fig.
-1(c).
-Considering the relatively smaller scale of land-covers in high-resolution remote sensing scenarios, we adopt two parallel convolution branches in LSM with the kernel size set as 3 and 1, respectively.
-Each branch is followed by a batch normalization layer with a ReLU6 activation function:
+Considering the relatively smaller scale of land-covers in high-resolution scenarios, we adopt two parallel convolution branches in LSM with the kernel size set as 3 and 1, respectively.
+Each branch is followed by a batch normalization layer with a ReLU6 activation:
 $\tilde{F^{\prime}}_{i}=\text{ReLU6}\left(\text{BatchNorm}\left(\text{Conv}_{i}\left(F_{MSD}\right)\right)\right)$ (4)
 where $i\in\left\{1,3\right\}$ denotes the kernel size, while $F_{MSD}$ denotes the output feature of MSD at the corresponding stage. The two branches are then merged by addition:
-$\tilde{F^{\prime}}=\tilde{F^{\prime}}_{1}\oplus\tilde{F^{\prime}}_{3}$ (5)
+$\tilde{F^{\prime}}=\tilde{F}_{1}^{{}^{\prime}}\oplus\tilde{F}_{3}^{{}^{\prime}}$ (5)
 Eventually, through a dropout layer, a 1 × 1 convolution consistent with the segmentation head in MSD, and an upsampling layer to the original image size, LSM obtains $\tilde{F}\in\mathbb{R}^{H\times W\times 3}$:
-$\tilde{F}=\text{Upsample}\left(\text{Conv}_{1}\left(\text{Drop}\left(\tilde{F^{\prime}}\right)\right)\right)$ (6)
+$\tilde{F}=\text{Upsample}\left(\text{Conv}_{1}\left(\text{Dropout}\left(\tilde{F^{\prime}}\right)\right)\right)$ (6)
 During the training process, we add LSM blocks with different input channels to the decoder at stage 2-4 and then attain the segmentation result by summing up, which is further used in calculating the auxiliary loss function.
 II-C Loss Functions
-To ensure the holistic performance of UNetMamba while achieving the supervisory role of LSM, loss functions utilized in training are divided into a principal one and an auxiliary one.
-Following prior works[3, 4, 5, 6, 7], we use a dice loss $L_{dice}$ and a cross-entropy loss $L_{ce}$ to jointly form our principal segmentation loss function:
+To ensure the holistic segmentation performance of UNetMamba, we select two classical loss functions to jointly form our principal loss:
 $L_{dice}=1-\frac{2}{N}\sum_{n=1}^{N}{\sum_{k=1}^{K}{\frac{\hat{y}_{k}^{\left(n\right)}y_{k}^{\left(n\right)}}{\hat{y}_{k}^{\left(n\right)}+y_{k}^{\left(n\right)}}}}$ (7)
 $L_{ce}=-\frac{1}{N}\sum_{n=1}^{N}{\sum_{k=1}^{K}{\hat{y}_{k}^{\left(n\right)}\log y_{k}^{\left(n\right)}}}$ (8)
-where $N$ denotes the number of samples while $K$ denotes the number of categories.
-$\hat{y}_{k}^{\left(n\right)}$ is the $k$-th element in one-hot encoding of the true semantic labels for sample $n$.
-And $y_{k}^{\left(n\right)}$ represents the confidence of sample $n$ belonging to category $k$.
-As to the auxiliary loss function, cross-entropy loss $L_{ce}$ is selected[7].
+where $L_{dice}$ and $L_{ce}$ represents dice loss and cross-entropy loss, respectively.
+And $N$ denotes the number of samples while $K$ denotes the number of categories.
+$\hat{y}_{k}^{\left(n\right)}$ is the $k$-th element in one-hot encoding of the true semantic labels for sample $n$, while $y_{k}^{\left(n\right)}$ is the confidence of sample $n$ belonging to category $k$.
+Moreover, to achieve the supervisory role of LSM, cross-entropy loss $L_{ce}$ is also selected as the auxiliary loss for local supervision.
 Therefore, the overall loss function is formulated as follows:
 $L=(L_{dice}+L_{ce})+\alpha L_{ce}$ (9)
-where the weight factor $\alpha$ is set as 0.4.
-Figure 2: Qualitative comparisons on the LoveDA dataset at size 1024 × 1024.
-(a) Origin, (b) GT, (c) our proposed UNetMamba, (d) BANet, (e) ABCNet, (f) MANet, (g) DC-Swin, (h) UNetFormer, and (i) RS3Mamba.
-Figure 3: Qualitative comparisons on the ISPRS Vaihingen dataset at size 1024 × 1024.
-(a) Origin, (b) GT, (c) our proposed UNetMamba, (d) BANet, (e) ABCNet, (f) MANet, (g) DC-Swin, (h) UNetFormer, and (i) RS3Mamba.
+where the weight factor $\alpha$ is set as 0.4 [7].
 III Experimental results
@@
 Moreover, the number of model parameters (Param), memory footprint (Memo) and floating point operation count (FLOPs) were selected for efficiency evaluation, while the mean F1-score (mF1), mean intersection over union (mIoU) and overall accuracy (OA) were chosen as accuracy evaluation metrics.
+Figure 2: Qualitative comparisons on the LoveDA dataset at size 1024 × 1024.
+(a) Origin, (b) GT, (c) our proposed UNetMamba, (d) BANet, (e) ABCNet, (f) MANet, (g) DC-Swin, (h) UNetFormer, and (i) RS3Mamba.
+Figure 3: Qualitative comparisons on the ISPRS Vaihingen dataset at size 1024 × 1024.
+(a) Origin, (b) GT, (c) our proposed UNetMamba, (d) BANet, (e) ABCNet, (f) MANet, (g) DC-Swin, (h) UNetFormer, and (i) RS3Mamba.
 III-C Performance Comparison
@@
 In background and agriculture, UNetMamba led the second place by 1.61% and 2.39%, respectively.
-This distinct advantage indicated that, thanks to MSD, our UNetMamba gained excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition.
+This distinct advantage indicated that, benefiting from MSD, our UNetMamba garnered excellent global perception ability, which made it capable of accurately segmenting large land-covers under such a high-resolution condition.
 The outstanding performance of UNetMamba in building, road and other small land-cover categories proved that the train-only LSM successfully enhanced the perception of local contents through supervision and generalization.
 The qualitative comparison results on LoveDA are illustrated in Fig.
-2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate.
+2, which intuitively shows that the semantic segmentation mask obtained by UNetMamba is more accurate with less omission and clearer boundaries.
 2) Comparison on the ISPRS Vaihingen Dataset: The proposed UNetMamba not only achieved SOTA performance on the challenging LoveDA dataset, but also performed noticeably well on the classic Vaihingen dataset.
 As shown in Tab.
-II, in terms of accuracy, our UNetMamba achieved improvements of 0.20% in mF1, 0.36% in mIoU, and 0.18% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) and low cost (225.71MB, 100.52G).
+II, in terms of accuracy, our UNetMamba achieved improvements of 0.20% in mF1, 0.36% in mIoU, and 0.18% in OA, while maintaining competitive efficiency through a lightweight model (14.76M) with low costs (225.71MB, 100.52G).
 Fig.
@@
 Extensive experiments conducted on two well-known remote sensing datasets demonstrate that UNetMamba not only outperforms other SOTA models, but also achieves light weight and low costs.
-In the future, we will continue to explore the potential of novel linear modeling mechanisms in efficient interpretation of high-resolution remote sensing images.
+In the future, we will continue to explore promising linear mechanisms to further improve the accuracy and efficiency of our UNetMamba.